OpenStudy (anonymous):

Angela is on her way from home in Jersey City into New York City for dinner. She walks 1 mile to the train station, takes the train of the way, and takes a taxi of the way to the restaurant. How far does Angela live from the restaurant?

OpenStudy (anonymous):

Could you please help me? Thanksalot.

OpenStudy (anonymous):

Angela is on her way from home in Jersey City into New York City for dinner. She walks 1 mile to the train station, takes the train 1/2 of the way, and takes a taxi 1/6 of the way to the restaurant. How far does Angela live from the restaurant?

OpenStudy (mayankdevnani):

This is the correct answer to that word problem: 1 + (4/5)x + (1/6)x = x which would become (29x + 30)/30 = x 29x + 30 = 30x 30x - 29x =30 x = 30 Therefore Angela lives 30 Miles from the restaurant About 42km

OpenStudy (mayankdevnani):

ok @aly12

OpenStudy (mayankdevnani):

sorry !!! answer is 30 km @aly12

OpenStudy (anonymous):

sir i have a question... ist 4/5? the given are, 1/2 ^ 1/6

OpenStudy (anonymous):

* &&

OpenStudy (mayankdevnani):

oh!!! sry

OpenStudy (mayankdevnani):

This is the correct answer to that word problem: 1 + (1/2)x + (1/6)x = x which would become

OpenStudy (mayankdevnani):

can you solve it?

OpenStudy (anonymous):

I'll try sir.... wait. :)

OpenStudy (mayankdevnani):

plzz

OpenStudy (anonymous):

why is my answer, 3miles? T_T Oh God. Im sorry, late reply/

OpenStudy (phi):

3 miles is correct: 1 + 3/2 + 3/6 adds up to 3

OpenStudy (anonymous):

@phi yuh u're right. its correct.

OpenStudy (anonymous):

its weird. im the one who's asking then suddenly some people wer asking me back. huhuu. newbie here. well anyway thank u so much for your help.:)

OpenStudy (phi):

You are the one that figured it out. I just checked the answer.

OpenStudy (anonymous):

this is really cool! Unbelievable/ Im starting to like this site. Well, still... thank u for checking. :******* :''>